OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Passing through (-5, -7) and (-8, -6)

OpenStudy (anonymous):

Change in Y/change in X equals gradient -6-(-7)/-8-(-5) =1/-3 1/-3=y-(-7)/x-(-5) 3y=-x-26 I THINK JUST DOUBLE CHECK

OpenStudy (anonymous):

math in biology!